A car driving 20m/s sees a deer in the road and brakes, slowing the car at a rate of -12 m/s^2.
How long does it take the car to stop?
If the deer is 10 meters from the car when it brakes, does the car hit it?
Would the deer survive if the acceleration rate had been -15 m/s^2?
Answer #1

= 20mle, a = - 120/82, ugao a V = + at & OF 20-12t 2 de vet + hat? = (20 X1.67) + (-12 x 1.622) - 16.7m > 10 To Yes, it will